OpenStudy (anonymous):

What is the equation of the line of symmetry for the graph of y = 3(x – 2)2 + 1? a. x = –2 b. x = 1 c. x = 2 d. x = 3

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

3(x – 2)^2 + 1 a(x-h)^2+k line of symmtery is x=h so in your case h=2 so x=2

OpenStudy (anonymous):

its 2 o r -2?

OpenStudy (anonymous):

what did I write?

OpenStudy (anonymous):

b/c dont u have to take ur sign with it? you for x = 2

OpenStudy (anonymous):

*wrote

OpenStudy (anonymous):

yes, in general case a(x-h)^2 symmerty line -> x=h a(x+h)^2 =a(x- -h)^2 --> x=-h

OpenStudy (anonymous):

so its a?

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

no, it is x=2

OpenStudy (anonymous):

buy why if before it has a -

OpenStudy (anonymous):

IF IT WERE 3(x + 2)^2 + 1 then it would be x=-2

OpenStudy (anonymous):

oh so its the opposite sign?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh i get it thank u!